OpenStudy (anonymous):

Please help. Write equations of the lines through the given point (a) parallel to the given line and (b) perpendicular to the given line. point (3,-2) line 5x-4y=8

OpenStudy (debbieg):

Solve 5x-4y=8 for y, so that you have that line equation in the form y=mx + b From that, you can find the slope m for the given line. 1. for the parallel line, what is the relationship between the slopes of parallel lines? Use that, and the slope you obtained from the given line, to determine the slope of the parallel line. Now you'll have the slope, and a point on the line, from which you can determine the equation of the line (either using point-slope form, or slope-intercept form, your choice really). 2. for the perdpendicular line, what is the relationship between the slopes of perpendicular lines? Use that, and the slope you obtained from the given line, to determine the slope of the perpendicular line. Proceed as in (1), but with the perpendicular slope.

OpenStudy (anonymous):

I got y=5/4x-2

OpenStudy (anonymous):

@DebbieG

OpenStudy (debbieg):

ok, so what's the slope?

OpenStudy (anonymous):

5/4

OpenStudy (anonymous):

for the parallel

OpenStudy (anonymous):

@DebbieG

OpenStudy (debbieg):

Sorry... yes. The slope of the given line is m=5/4, and parallel lines have equal slope. So now the problem is reduced to: find the equation for the line through (3,-2) with m=5/4

OpenStudy (anonymous):

I got y=5/4x+5 for a

OpenStudy (anonymous):

@DebbieG